# 2]organization=ASML, city=Veldhoven, country=The Netherlands

### v1

@@ 5.2.2 Parameter mappings for proportionally-damped mechanical systems
In summary, the above derivations provide explicit parameter mappings and their Jacobians for both general-viscously and proportionally damped mechanical systems. Together with the Gauss–Newton scheme in (45), these expressions enable the structured identification of modal models from the estimated additive parameters, thereby completing the solution of the IPEM problem (22).
@@ 6 Experimental validation
Next, the two-stage identification procedure is experimentally validated on the experimental wafer-stage system introduced in Section 2, thereby providing contribution C3.
@@ 6.2 Model structure and model-order selection
Next, the model structure and corresponding model orders used for parametric identification are specified. The general damping model structure is considered; however, it is noted that since the wafer-stage system is well described by a proportional damping assumption, the general formulation yields only a slight improvement in performance for this specific case. The wafer-stage system exhibits $n_{\mathrm{rbm}}=3$ rigid-body modes in the out-of-plane direction, corresponding to translation along the $z$-axis and rotations around the $x$- and $y$-axes. The number of flexible modes $n_{\mathrm{flex}}$ is determined using the CMIF Shih1988ComplexEstimation , which is computed from the squared singular values of the FRF matrix at each frequency. The corresponding modal frequencies are identified from the peak locations in the CMIF, while mode multiplicity is inferred from the number of singular values peaking at the same frequency. From the CMIF of the FRF dataset, shown in Figure 4, $n_{\mathrm{flex}}=17$ distinct flexible modes are identified. Thus, the used model structure consists of $n_{\mathrm{rbm}}=3$ rigid-body modes, $n_{\mathrm{flex}}=17$ flexible modes, and an additional DC gain to capture the effect of higher-order dynamics, which results in a 40th-order parametric model.
Figure 4: CMIF plot used for model order selection and initialization of the mode frequencies, with $\sigma_{i}(\omega_{k})$ denoting the $i$th singular value of the measured FRF dataset at frequency $k$. The frequency locations of the flexible modes are indicated by peaks in the singular-value plot, with multiplicity determined by the number of singular values peaking at each frequency.
@@ 6.4 Initialization
Due to the non-convexity of the cost function in (21), accurate initialization of the estimator (40) is essential. To initialize the stage one estimation problem, initial pole locations must be provided, which enable the numerators to be computed using the convex LS problem (23). The pole locations are constructed from initial estimates $\omega_{i}$ and $\zeta_{i}$, where the natural frequencies $\omega_{i}$ are set to the CMIF peak frequencies, and damping ratios $\zeta_{i}$ are uniformly assigned a value of 0.01, representative of lightly damped flexible modes. It is noted that subsequent RIV estimation demonstrated a low sensitivity to the initial choice of damping coefficient, which may alternatively be computed using nonparametric modal parameter estimation techniques DeKraker2004ADynamics or FEM modeling.
@@
Figure 8: Bode magnitude plot (top) with the corresponding residual (bottom) for single plant entries. Shown are the FRF measurement $\mathbf{G}(\omega_{k})$ ( ) and the estimated parametric modal model $\hat{\mathbf{P}}_{m}(j\omega_{k})$ ( ).
Figure 9: Estimated mode shapes of the experimental wafer-stage system (actuator side) for the first 12 modes. The shapes are interpolated using the thin-plate spline method introduced in Voorhoeve2021IdentifyingStage . The first 3 modes correspond to rigid-body modes, while the remaining 9 are flexible modes.
@@ 6.5 Results
The two-stage parametric identification procedure is applied to estimate a parametric model of the wafer-stage system. In the first stage, an additive model is estimated using 10 iterations of (40), and in the second stage this solution is used to estimate the modal model over 40 iterations of (45). In Figure 5, the evolution of the cost function (21) is shown. It is emphasized that the RIV method does not guarantee monotonic convergence in general but typically demonstrates fast convergence to a locally optimal solution. Figure 6 shows the singular value distribution of the estimated residue matrices after the first stage ($j=10$). This distribution reflects the extent to which the model approximates an ideal modal form. In an exact modal representation of flexible modes, the residue matrices in (12) must be exactly rank one. However, the estimated residue matrices contain additional nonzero singular values, with the largest discrepancy for the modes around 500 Hz and 1800 Hz, motivating the need for further refinement. The rigid body mode shows three dominant singular values, consistent with $n_{\mathrm{rbm}}=3$, and since the DC gain must approximate many higher‐order modes, its residue matrix is expected to be full rank.
Next, the second-stage IPEM problem defined in (45) is solved. First, an initial modal model is computed using the SVD-based decomposition, which is subsequently refined through GN iterations. The frequency response of the identified modal model is shown in Figure 8, together with the FRF measurement used in the estimation. Furthermore, Figure 8 shows the frequency response of individual plant entries, along with the corresponding model error. These figures shows a good agreement between the modal model and the FRF over the complete frequency range. In the frequency range around 500 Hz, the modeling error is slightly larger, suggesting either the presence of higher mode multiplicities or of unmodeled dynamics not captured by the modal-model structure.
Finally, Figure 9 shows the estimated mode shape vectors of the experimental wafer-stage system at the actuator side. In this case, the real mode shapes corresponding to the proportional damping model structure are presented, which are interpolated using thin-plate splines (see Voorhoeve2021IdentifyingStage for details). The first three modes correspond to rigid-body motion, while the remaining modes represent flexible dynamics. These mode shapes provide insight into the spatial deformation patterns of the wafer-stage system, which can be used in both controller and system design.
@@ 7 Conclusions
This paper introduced a parametric identification method for complex, multivariable mechanical systems in a modal model structure, that accommodates both general‐viscous and proportional damping formulations. By decomposing the structured identification problem into two sequential estimation stages, the proposed framework enables computationally efficient estimation of modal models for high‐dimensional systems. Experimental validation on a prototype wafer‐stage system with 13 actuators and 4 sensors produced an accurate 40th‐order modal model, comprising three rigid‐body modes and 17 flexible modes, that closely matches the measured FRF. These results demonstrate the method’s effectiveness and robustness in modeling the dynamics of high‐dimensional industrial systems.
Appendix A Derivation of the modal forms
@@ A.1 General-viscously damped systems
$(\mathbf{A}-\boldsymbol{\lambda}_{i}\mathbf{E})\mathbf{v}_{i}=\mathbf{0},\quad i=\{1\dots 2n\},$ (57)
where $\mathbf{v}_{i}\in\mathbb{C}^{2n}$ is the complex mode-shape vector, normalized according to $\mathbf{v}_{i}^{\top}\mathbf{E}\mathbf{v}_{i}=1$, and $\boldsymbol{\lambda}_{i}\in\mathbb{C}$ is the corresponding eigenvalue given in (5). The state relation described by the matrices (56) enables the modal matrix $\mathbf{V}=[\mathbf{v}_{1},\ldots,\mathbf{v}_{2n}]$ to be partitioned as $\mathbf{V}=\left[\mathbf{U}^{\top},\boldsymbol{\Lambda}\mathbf{U}^{\top}\right]^{\top}$, with $\mathbf{U}\in\mathbb{C}^{n\times 2n}$ containing the unique mode shape information, and the matrix $\boldsymbol{\Lambda}=\operatorname{diag}([\boldsymbol{\lambda}_{1},\ldots,\boldsymbol{\lambda}_{2n}])$ contains the eigenvalues along the main diagonal. The descriptor system (55) is transformed to a modal system with states $\mathbf{x}_{m}\in\mathbb{R}^{2n}$ by substituting $\mathbf{x}(t)=\mathbf{V}\mathbf{x}_{m}(t)$ followed by premultiplication with $\mathbf{V}^{\top}$ to obtain
@@
$\mathbf{P}_{m}(s)$ $=\sum_{i=1}^{2n}\frac{\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}_{r,i}^{\top}}{s-\boldsymbol{\lambda}_{i}}.$ (60)
For real-valued state matrices in (56), the eigenvalues $\boldsymbol{\lambda}_{i}$ and mode-shape vectors $\mathbf{v}_{i}$ appear as complex conjugate pairs. This property enables the first-order form to be partitioned according to
$\mathbf{P}_{m}(s)$ $=\sum_{i=1}^{n}\frac{\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}_{r,i}^{\top}}{s-\boldsymbol{\lambda}_{i}}+\frac{\bar{\boldsymbol{\psi}}_{l,i}\bar{\boldsymbol{\psi}}_{r,i}^{\top}}{s-\bar{\boldsymbol{\lambda}}_{i}},$ (61)
@@ A.3 State‐space representation
where $\boldsymbol{\Lambda}_{n}=\operatorname{diag}([\boldsymbol{\lambda}_{1},\ldots,\boldsymbol{\lambda}_{n}])$ contains the $n$ unique eigenvalues along the diagonal. Applying the similarity transform yields the transformed system

### v2

@@ 5.2.2 Parameter mappings for proportionally damped mechanical systems
5.3 Summary
The above derivations provide explicit parameter mappings and their associated Jacobians for both generally viscously damped and proportionally damped mechanical systems. These parameter mappings and their Jacobians are used to formulate and solve the IPEM problem, enabling the structured identification of modal models from the additive parameters estimated in the first stage and thereby completing the solution to the IPEM problem (22). A complete overview of the proposed two-stage MIMO modal identification framework is provided in Procedure 1.
Algorithm 1 Two-stage approach for MIMO modal model estimation
1:FRF dataset $\{G(j\omega_{k})\}^{N}_{k=1}$, initialization $\{\omega^{\langle 0\rangle}_{i},\zeta^{\langle 0\rangle}_{i}\}^{n_{\mathrm{flex}}}_{i=1}$, RIV tolerance $\epsilon_{1}$, IPEM tolerance $\epsilon_{2}$
2:Stage 1: additive model estimation
3:Construct $\bar{A}_{i}(s)$ from $\{\omega^{\langle 0\rangle}_{i},\zeta^{\langle 0\rangle}_{i}\}^{n_{\mathrm{flex}}}_{i=1}$.
4:Solve the linear LS problem (23) to obtain $\hat{\boldsymbol{\eta}}$ and construct $\boldsymbol{\beta}^{\langle 0\rangle}$.
5:Solve the non-linear LS problem (21) using the RIV Algorithm 1 to obtain $\hat{\boldsymbol{\beta}}$.
6:Compute $\hat{\boldsymbol{\Sigma}}_{\hat{\boldsymbol{\beta}}}$ using (41)
7:Stage 2: projection to modal model
8:Construct $\boldsymbol{\rho}^{\langle 0\rangle}$ from $\hat{\boldsymbol{\beta}}$ using SVDs
9:Solve the IPEM problem (22) using Algorithm 2 to obtain $\hat{\boldsymbol{\rho}}$.
10:Modal parameter vector estimate $\hat{\boldsymbol{\rho}}$
@@ 6 Experimental validation
Next, the two-stage identification procedure is experimentally validated on an experimental wafer-stage system. First, the experimental setup is introduced, followed by a discussion of the estimation process, and finally the results are presented.
6.1 Experimental setup: experimental wafer-stage system
The considered experimental setup depicted in Figure 2 is a prototype wafer stage system, which is designed as a lightweight mechanical structure to enable fast motions with high accelerations. The stage is magnetically levitated by passive gravity compensators positioned at its corners, establishing a mid-air equilibrium, and is actively controlled in six motion degrees of freedom at a sampling rate of 10 kHz, achieving sub-micrometer accuracy. The system contains 17 actuators, with 13 voice-coil actuators in the $z$-direction and two Lorentz actuators for each $x$- and $y$-directions. To measure the position and orientation of the stage, the system includes seven positioning sensors, consisting of four linear encoders for the $z$-direction, two capacitance sensors for the $x$-direction, and a single capacitance sensor for the $y$-direction.
The complete input–output dynamics of the wafer-stage system are described by a $7\times 17$ transfer matrix. To simplify the presentation of the results, only the measurements in the out-of-plane direction are considered, namely translation along the $z$ axis and rotations around the $x$ and $y$ axes. The corresponding input–output dynamics of this subset of actuators and sensors are described by a $4\times 13$ transfer matrix. A schematic overview of the actuator and sensor layout in the out-of-plane direction is provided in Figure 2. In addition, in Figure 4 and Figure 4 a detailed overview is provided of the experimental setup. The lightweight design of the stage results in high-order, lightly damped flexible dynamics that limit system performance, making accurate parametric models essential for achieving further performance improvements.
@@ 6.2 Nonparametric modeling
Figure 3: Overview of the experimental wafer-stage setup. The sensory equipment is mounted on the metrology frame (A), which is isolated from the baseframe (B) through an air suspension system (C). The linear encoders for the out-of-plane measurements and the capacitance sensors for the in-plane measurements are mounted on the pillars (D), which are attached to the metrology frame. The actuation system is mounted on the force frame (E), and the stage (F) is isolated from the force frame by a set of gravity compensators.
Figure 4: Overview of the actuation system. The force frame is shown extracted from the machine, with the locations of the in-plane Lorentz motors for the $x$-direction shown in (A) and the actuator locations for the $y$-direction shown in (B). For the out-of-plane motion direction, 13 voice-coil actuators are used, with a single actuator shown in (C). Four passive gravity compensators are positioned at the corners (D) to compensate the gravitational load of the stage.
@@ 6.3 Model structure and model-order selection
Next, the model structure and corresponding model orders used for parametric identification are specified. The general damping model structure is considered; however, it is noted that since the wafer-stage system is well described by a proportional damping assumption, the general formulation yields only a slight improvement in performance for this specific case. The wafer-stage system exhibits $n_{\mathrm{rbm}}=3$ rigid-body modes in the out-of-plane direction, corresponding to translation along the $z$-axis and rotations around the $x$- and $y$-axes.
The number of flexible modes $n_{\mathrm{flex}}$ is determined using the CMIF Shih1988ComplexEstimation , which is computed from the squared singular values of the FRF matrix at each frequency. The corresponding modal frequencies are identified from the peak locations in the CMIF, while mode multiplicity is inferred from the number of singular values peaking at the same frequency. From the CMIF, shown in Figure 5, $n_{\mathrm{flex}}=17$ distinct flexible modes are identified. Thus, the used model structure consists of $n_{\mathrm{rbm}}=3$ rigid-body modes, $n_{\mathrm{flex}}=17$ flexible modes, and an additional DC gain to capture the effect of higher-order dynamics, which results in a 40th-order parametric model.
Figure 5: CMIF plot used for model order selection and initialization of the mode frequencies (left) with additional zoomed-in view (right), with $\sigma_{i}(\omega_{k})$ denoting the $i$th singular value of the measured FRF dataset at frequency $k$. The frequency locations of the flexible modes are indicated by peaks in the singular-value plot, with multiplicity determined by the number of singular values peaking at each frequency.
@@ 6.5 Initialization
Due to the non-convexity of the cost function in (21), accurate initialization of the estimator (40) is essential. To initialize the stage one estimation problem, initial pole locations must be provided, which enable the numerators to be computed using the convex LS problem (23). The pole locations are constructed from initial estimates $\omega_{i}$ and $\zeta_{i}$, where the natural frequencies $\omega_{i}$ are set to the CMIF peak frequencies, and damping ratios $\zeta_{i}$ are uniformly assigned a value of 0.01, representative of lightly damped flexible modes.
@@
Figure 9: Bode magnitude plot (top) with the corresponding residual (bottom). Shown are the FRF measurement $\mathbf{G}(\omega_{k})$ ( ), the estimated parametric modal model $\hat{\mathbf{P}}_{m}(j\omega_{k})$ ( ), and the residual ${\mathbf{E}}_{m}(j\omega_{k})$ ( ).
Figure 10: Experimentally identified mode shapes of the experimental wafer-stage system (actuator side) for the first 12 modes. The shapes are interpolated using the thin-plate spline method introduced in Voorhoeve2021IdentifyingStage . The first 3 modes correspond to rigid-body modes, while the remaining 9 are flexible modes.
Figure 11: Theoretical mode shapes obtained from the FEM model of the wafer-stage system for the first 12 nontrivial out-of-plane modes. The shapes are interpolated onto a regular grid using the same thin-plate spline procedure as used for the experimentally identified modes. The first 3 modes represent rigid-body motion, while the remaining 9 correspond to flexible structural modes.
@@ 6.6 Results
The two-stage parametric identification procedure is applied to estimate a parametric model of the wafer-stage system. In the first stage, an additive model is estimated by solving (21) using the RIV approach. In the second stage, the resulting estimate is projected onto a modal model by solving the IPEM problem (22), which is implemented using available gradient-based numerical solvers in MATLAB. In Figure 7, the evolution of the cost function (21) is shown. It is emphasized that the RIV method does not guarantee monotonic convergence in general but typically demonstrates fast convergence to a locally optimal solution. Figure 7 shows the singular value distribution of the estimated residue matrices after the first stage ($j=10$). This distribution reflects the extent to which the model approximates an ideal modal form. In an exact modal representation of flexible modes, the residue matrices in (12) must be exactly rank one. However, the estimated residue matrices contain additional nonzero singular values, with the largest discrepancy for the modes around 500 Hz and 1800 Hz, motivating the need for further refinement. The rigid body mode shows three dominant singular values, consistent with $n_{\mathrm{rbm}}=3$, and since the DC gain must approximate many higher‐order modes, its residue matrix is expected to be full rank.
Next, the second-stage IPEM problem defined in (45) is solved. First, an initial modal model is computed using the SVD-based decomposition, which is subsequently refined through the projection step. The frequency response of the identified modal model is shown in Figure 9, together with the FRF measurement used in the estimation. Furthermore, Figure 9 shows the frequency response of individual plant entries, along with the corresponding model error. These figures shows a good agreement between the modal model and the FRF over the complete frequency range. In the frequency range around 500 Hz, the modeling error is slightly larger, suggesting either the presence of higher mode multiplicities or of unmodeled dynamics not captured by the modal-model structure.
Finally, Figure 11 visualizes the estimated mode-shape vectors at the actuator side. Real-mode shape-vectors are provided corresponding to the proportionally damped model structure, which are interpolated using thin-plate splines (see Voorhoeve2021IdentifyingStage for details). In addition, Figure 11 shows the theoretical mode shapes obtained from a Finite-Element Method (FEM) analysis of the wafer-stage system, enabling a qualitative comparison between the estimated and modeled dynamics. Details of the FEM model are provided in Appendix B. The first three modes correspond to rigid-body motion, while the remaining modes represent flexible dynamics. The first nine flexible modes show a clear correspondence between the experimental and FEM mode shapes, whereas the higher modes show deviations due to the limited spatial resolution of the available actuators and sensors. These mode shapes provide valuable insight into the spatial deformation behavior of the wafer-stage system and support both controller design and system analysis.
@@ 7 Conclusions
This paper introduced a parametric identification method for complex, MIMO mechanical systems in a modal model structure, that accommodates both general‐viscous and proportional damping formulations. By decomposing the structured identification problem into two sequential estimation stages, the proposed framework enables computationally efficient estimation of modal models for high‐dimensional systems. Experimental validation on a prototype wafer‐stage system with 13 actuators and 4 sensors produced an accurate 40th‐order modal model, comprising three rigid‐body modes and 17 flexible modes, that closely matches the measured FRF. These results demonstrate the method’s effectiveness and robustness in modeling the dynamics of high‐dimensional industrial systems.
Appendix A Modal representations of mechanical systems
@@ A.1 General-viscously damped systems
$(\mathbf{A}-\lambda_{i}\mathbf{E})\mathbf{v}_{i}=\mathbf{0},\quad i=\{1\dots 2n\},$ (57)
where $\mathbf{v}_{i}\in\mathbb{C}^{2n}$ is the complex mode-shape vector, normalized according to $\mathbf{v}_{i}^{\top}\mathbf{E}\mathbf{v}_{i}=1$, and $\lambda_{i}\in\mathbb{C}$ is the corresponding eigenvalue given in (5). The state relation described by the matrices (56) enables the modal matrix $\mathbf{V}=[\mathbf{v}_{1},\ldots,\mathbf{v}_{2n}]$ to be partitioned as $\mathbf{V}=\left[\mathbf{U}^{\top},\boldsymbol{\Lambda}\mathbf{U}^{\top}\right]^{\top}$, with $\mathbf{U}\in\mathbb{C}^{n\times 2n}$ containing the unique mode shape information, and the matrix $\boldsymbol{\Lambda}=\operatorname{diag}([\lambda_{1},\ldots,\lambda_{2n}])$ contains the eigenvalues along the main diagonal. The descriptor system (55) is transformed to a modal system with states $\mathbf{x}_{m}\in\mathbb{R}^{2n}$ by substituting $\mathbf{x}(t)=\mathbf{V}\mathbf{x}_{m}(t)$ followed by premultiplication with $\mathbf{V}^{\top}$ to obtain
@@
$\mathbf{P}_{m}(s)$ $=\sum_{i=1}^{2n}\frac{\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}_{r,i}^{\top}}{s-\lambda_{i}}.$ (60)
For real-valued state matrices in (56), the eigenvalues $\lambda_{i}$ and mode-shape vectors $\mathbf{v}_{i}$ appear as complex conjugate pairs. This property enables the first-order form to be partitioned according to
$\mathbf{P}_{m}(s)$ $=\sum_{i=1}^{n}\frac{\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}_{r,i}^{\top}}{s-\lambda_{i}}+\frac{\bar{\boldsymbol{\psi}}_{l,i}\bar{\boldsymbol{\psi}}_{r,i}^{\top}}{s-\bar{\lambda}_{i}},$ (61)
@@ A.3 State‐space representation
where $\boldsymbol{\Lambda}_{n}=\operatorname{diag}([\lambda_{1},\ldots,\lambda_{n}])$ contains the $n$ unique eigenvalues along the diagonal. Applying the similarity transform yields the transformed system
@@
Appendix B FEM model OAT
The FEM model is derived from a Siemens NX geometry of the wafer stage and discretized using tetrahedral elements, resulting in a mesh with 85 262 nodes and 259 185 elements. The wafer stage is modeled as a thin rectangular structure with a width of 0.6 [m], length 0.6 [m] and a height of 45 [mm], which for the considered material properties leads to a mass of approximately 8.6 [kg]. The FEM model is used to compute the theoretical mode shapes and eigenfrequencies of the experimental wafer-stage system, which serve as a reference for validating the experimentally identified modes. The FEM model can be seen in Figure 12.
Figure 12: FEM model of the wafer-stage system.